# FlashBack : Efficient Retrieval-Augmented Language Modeling for Long Context Inference

### v3

@@ Abstract
Retrieval-Augmented Language Modeling (RALM) by integrating large language models (LLM) with relevant documents from an external corpus is a proven method for enabling the LLM to generate information beyond the scope of its pre-training corpus. Previous work utilizing retrieved content by simply prepending it to the input poses a high runtime issue, which degrades the inference efficiency of the LLMs because they fail to use the Key-Value (KV) cache efficiently. In this paper, we propose FlashBack, a modular RALM designed to improve the inference efficiency of RALM with appending context pattern while maintaining decent performance after fine-tuning by Low-Rank Adaption. FlashBack appends retrieved documents at the end of the context for efficiently utilizing the KV cache instead of prepending them. And we introduce Marking Token as two special prompt tokens for marking the boundary of the appending context during fine-tuning. Our experiments on testing generation quality show that FlashBack can remain decent generation quality in perplexity. And the inference speed of FlashBack is up to 4×4$\times$4 × faster than the prepending counterpart on a 7B LLM (Llama 2) in the runtime test. Via bypassing unnecessary re-computation, it demonstrates an advancement by achieving significantly faster inference speed, and this heightened efficiency will substantially reduce inferential cost.
@@ 1 Introduction
Large language models (LLMs) based on the Transformer architecture (Vaswani et al., 2023) such as GPT, Llama and OPT, etc (Brown et al., 2020; Touvron et al., 2023; Zhang et al., 2022) require enormous computational resources to keep their knowledge updated (Meng et al., 2023). Thus, Retrieval-Augmented Language Modeling (RALM) has emerged as a popular approach, enabling content generation that leverages external corpora to extend beyond the knowledge inherent in the model’s parameters, thereby reducing the computational cost of capturing up-to-date knowledge. The effectiveness of using RALM to enhance LLMs’ performance has been proven by previous studies(Guu et al., 2020; Izacard et al., 2022; Lewis et al., 2021; Wang et al., 2023b).
Previous studies build RALM through the pre-training of LLM(Borgeaud et al., 2022), or engage in the continual pre-training of existing RALMs (Izacard et al., 2022; Wang et al., 2023a). Recent works utilizing retrieved content by simply prepending retrieved contents to the input without updating the language models exhibit the impressive potential of adapting off-the-shelf LLMs to the retrieved content (Ram et al., 2023; Shi et al., 2023).
However, these works are introduced with limitations. First, the off-the-shelf LLMs are not inherently trained to incorporate retrieved content, and extensive pre-training of LLMs for building RALM incurs high computational costs (Lin et al., 2023). Second, although in-context methods have been effectively applied on off-the-shelf LLMs (Ram et al., 2023; Shi et al., 2023), recent research indicates that the bottleneck of these methods is redundancy and inefficiency (Asai et al., 2024). And our work analyze that prepending context pattern they used poses high computational cost in long context scenario, which tremendously increases the inference time (see Figure 1). To be specific, retrieved content changes with every retrieval stride tokens, and the computed key-value (KV) cache of the entire context is discarded, requiring re-computation for each change. Text generation process with long context can substantially increase the inference runtime, and the attention KV cache grows even larger on LLM with hundreds of billions of parameters(Pope et al., 2022). Under these circumstances, the re-computation of the KV cache becomes a computationally intensive task, entailing significant costs for In-Context RALM(Ram et al., 2023).
@@
Figure 1: A series of inference tests with simulated input and retrieved content on OPT and Llama 2 for comparing Appending Pattern (ours) with Prepending Pattern. (Maximum sequence length is denoted under the model name)
@@
In this paper, we propose FlashBack to overcome these issues we mentioned above, a method that fine-tuning off-the-shelf LLMs efficiently through adaption without further pre-training of LLMs as shown in figure 2, and to address the inefficiency issue by breaking the dependent of the input to retrieved content in the context. To be specific, instead of prepending the retrieved content directly to the input, we propose a new context pattern, Appending Context Pattern, in which the retrieved content is appended at the end of the input (see figure 3). In our method, the KV cache of each input is static during inference, and the re-computation of the KV cache takes place only in the part of the retrieved content. Although there are advantages to applying this new context pattern, it also presents an issue that degrading the performance of the RALM. Our appending pattern breaks the semantic coherence of the context, which leads to a loss in perplexity when applying it to existing RALM. To solve this issue, inspired by (Ren et al., 2023), FlashBack introduces tunable Marking Token to adapt to our appending context pattern for restoring performance, which is different from pre-training the entire model weights.
@@
Notably, we use LoRA (Hu et al., 2021) to fine-tune the Marking Token while keeping both the retriever and the LLM frozen. Therefore, FlashBack is a plug-and-play approach that can cooperate with other retrieval-oriented methods aiming for efficient inference and context-model alignment. When using an LLM such as OPT-6.7B as a reader, user can run the whole token-tuning process with a single 24GB-RAM GPU in BF16 precision.
@@
  • We find the Appending Context Pattern can improve the inference efficiency of In-Context RALMs. (Up to 4×4$\times$4 × faster in 7B Llama 2 as shown in figure 1)
  • We exploit Marking Token that can be used for adapting the LLMs to unseen context pattern by fine-tuning with LoRA. And this alignment method with LoRA can keep the LLM intact and avoid heavily damaging the knowledge integrity of the LLM.
  • Our modeling, the FlashBack, is orthogonal to other methods using off-the-shelf LLM. It has considerable potential for enhancement combined with other methods in future works.
Figure 2: An overview of the FlashBack pipelines. The left side of the diagram is the fine-tuning phase of our model with Marking Token and LoRA module. The inference procedure is on the right.
Figure 3: An illustration of the Appending Context Pattern(ours) comparing with Prepending Context Pattern.
@@ 2 Background and Related Work
2.1 Retrieval-Augmented Language Models
$k$NN-LM is a pioneering method that substantiated its capability as RALM, and it suggests that learning similarity functions between contexts may be a better solution than predicting the next word (Khandelwal et al., 2020). However, a recent investigation (Wang et al., 2023c) has revealed that the perplexity of $k$NN-LM exhibits improvement for a limited set of tokens but exacerbates predictions for the majority of tokens, particularly when generating lengthy sequences. As a result, this characteristic adversely affects the overall quality of text generation when compared to GPT-2 (Wang et al., 2023c). RALM also can be based on models using encoder-decoder structure(Huang et al., 2023; Lewis et al., 2021) and Atlas (Izacard et al., 2022) building upon the T5 language model (Raffel et al., 2023) stands out as a state-of-art RALM. Other RALMs (Lin et al., 2023; Ram et al., 2023; Shi et al., 2023; Xu et al., 2023) studied for utilizing decoder-only language models such as GPT, Llama and OPT, etc(Brown et al., 2020; Touvron et al., 2023; Zhang et al., 2022) or building their own autoregressive RALM (Borgeaud et al., 2022).
2.2 Retrieve-Read RALM
Previous works (Borgeaud et al., 2022; Lin et al., 2023; Ram et al., 2023; Shi et al., 2023) have created distinct modules for document selection and document reading. In the recent RALM framework, particularly for those employing LLMs, the imperative is to align retrieved documents with the specific requirements of the LLMs (Gao et al., 2024). AAR, REPLUG, and UPRISE are fine-tuning retrievers with frozen LLMs to achieve the alignment (Cheng et al., 2023; Shi et al., 2023; Yu et al., 2023). RETRO uses frozen retrievers to build their own RALM (Borgeaud et al., 2022). In-Context RALM uses a frozen retriever for document selection and a frozen LLM for document reading without undergoing additional training for either the LLM or the retriever (Ram et al., 2023). RAVEN (Huang et al., 2023) continually pre-train ATLAS (Izacard et al., 2022) for improving its in-context capabilities, which aligns the model to specific prompting strategy, and they also proposed Fusion-in-Context Learning for incorporating more in-context examples during its inference process.
Hence, our modeling is close to the genre of RALM that uses decoder-only LLMs (Auto-regressive Models) and Retrieval-Read RALM that uses distinct modules for document selection and document reading (Borgeaud et al., 2022; Ram et al., 2023; Lin et al., 2023; Shi et al., 2023). Also, our implementation on input context is classified as Input augmentation by recent study(Asai et al., 2024). After all, This type of RALM has recently been defined as Modular RAG (Retrieval-Based-Generation), and it is increasingly becoming the dominant norm in the RAG domain (Gao et al., 2024).
@@ 3.1 RALM with In-Context-Learning
For the trade-off between runtime cost and model performance, retrieval stride $s$ and query length $\ell$ are proposed for less frequently performing the retriever call (at least, every s > 1 token). Therefore, the retrieval would be performed for every $s$ tokens, using the last $\ell$ tokens as designated retrieved content for input.
@@
$p(x_{1},...,x_{n})=\prod_{j=0}^{n_{s}-1}\prod_{i=1}^{s}p_{\theta}(x_{s\cdot j+i}|[\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)});x_{<s\cdot j+i}]),$ (2)
@@
Our work followed this mechanism of retrieval besides using different context patterns because this mechanism uses distinct retrieve modules and readers (LLMs) so that they can be optimized independently.
@@ 3.2 Context pattern
In the decoder-only transformer-based models, the computation of attention modules is related to the query of the current token and the key-value representations of preceding tokens. (Vaswani et al., 2023). During inference, the key-value of the current token would be saved in the KV cache, eliminating the requirement in the subsequent step. We find that prepending retrieved content to the input has been a prevalent in previous methods(Ram et al., 2023; Shi et al., 2023). However, the key-value computed for previous tokens becomes obsolete, and a re-computation of new key-value is required for each retrieval step since the retrieved content varies with each retrieval.
To avoid the extensive cost of re-computation, we remove the auto-regressive dependency between retrieved content and input in context by using an appending context pattern. Specifically, we append retrieved content to input (at the end of the input) as shown in Figure 4.
@@
$p(x_{1},...,x_{n})=\prod_{j=0}^{n_{s}-1}\prod_{i=1}^{s}p_{\theta}(x_{s\cdot j+i}|[x_{<s\cdot j+i};\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)})]).$ (3)
In this case, for each input, the change of retrieved content $\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)})$ would not require a re-computation of new value therefore preserving the computed value of previous tokens in KV cache $x_{<s\cdot j+i}$. (Algorithm pseudo code in Appendix A.)
@@ 3.4 Marking Token and Fine-tuning Choice
Since LLMs are not aligned explicitly to our appending pattern, we use Marking Token and LoRA techniques to adapt them to the appending pattern while keeping origin model weights frozen so that the alignment is achieved without modifying the inherent ability of the LLMs. Marking Tokens are represented as two special prompt tokens denoting <MARK_L>,<MARK_R> which are added into the vocabulary of the model, and they are used for marking the boundary of the retrieved content in context (see Figure 2). Then, we fine-tune the model using Low-Rank Adaption (LoRA) (Hu et al., 2021) with Marking Tokens applied for adapting the appending pattern. LoRA is a parameter-efficient fine-tuning (PEFT) for fine-tuning a subset of parameters in the pre-trained model while targeting for obtaining comparable performance to full model fine-tuning. Recent study suggests that forgetting is inevitable in fine-tuning but PEFT method can facilitate less forgetting during fine-tuning process which does not heavily damage the knowledge integrity of the pre-trained LLM (Kalajdzievski, 2024). In FlashBack, we opt for LoRA with the aim of demonstrating how fine-tuning a relatively minimal set of model weights enables our model to adapt to the appending context pattern while keeping pre-training model weights frozen. Our experiments show that FlashBack has significantly boosted the inference speed, and it can still maintain competitive perplexity (PPL) after fine-tuning. During fine-tuning, all parameters are frozen except for the embedding of <MARK_L>, <MARL_R>, and LoRA modules applied to attention layers.
@@ 3.5 FLOPs analysis of Marking Token and LoRA
If the rank of weight matrices of LoRA is $r$, then the projection matrix dimensions are $[h,r]$. The average length of retrieved content is $d$. We only equip LoRA modules, two projection matrices, for Key and Value matrices. Overall, the FLOPs of LoRA modules in appending context pattern can be formulated as: [^1] We only consider multiplication operations and omit some addition operations.
$\begin{split}C_{1}&=2l(4bThr+bTh)+2l\sum_{i=1}^{T/s}\left[4b(d+s)hr+b(d+s)h\right]\\ &=\frac{2l(4r+1)bhT(d+2s)}{s}\end{split}$ (5)
@@
$C_{\text{decreament}}=C_{0}-C_{1}=\frac{2lTbh\left[(T+s)h-(4r+1)(d+2s)\right]}{s}$ (6)
@@ 3.6 Retriever
Our experiment used a sparse model, the BM25 (Robertson and Zaragoza, 2009), for demonstrating our idea. In practice, any other retrievers can also be used for the document selection process, and FlashBack intrinsically supports switching to other retrievers in a plug-and-play manner.
@@ 4.1 Run-time Improvement
We test OPT, GPT-2, and Llama2 in different model sizes and employ simulated retrieved content and input in inference speed tests. Our experiment encompassed tests under varying lengths of input. Additionally, it involved examining the inference time under two distinct context patterns.

### v4

# FlashBack : Efficient Retrieval-Augmented Language Modeling for Fast Inference

@@ Abstract
Retrieval-Augmented Language Modeling (RALM) by integrating large language models (LLM) with relevant documents from an external corpus is a proven methodology for enabling the LLM to generate information beyond the scope of its pre-training corpus. Previous work by retrieving a set of tokens iteratively with retrieved content prepending to the input poses a high run-time issue, which degrades the inference efficiency of the LLMs because they fail to use the Key-Value (KV) cache efficiently. We propose FlashBack, a modular RALM designed to improve the inference efficiency of RALM with the appending context pattern while maintaining decent performance after fine-tuning by Low-Rank Adaption. FlashBack appends retrieved documents at the end of the context to efficiently utilize the KV cache. We also introduce the Marking Token as two special prompt tokens for marking the appending context during fine-tuning. Our experiments show that FlashBack can improve language modeling performance in the perplexity metric. We proved that the Marking Token is a usable add-on when fine-tuning models on specific context patterns. By bypassing unnecessary recomputation, FlashBack achieves fast inference speed with long context input. The inference speed is up to 4×4$\times$4 × faster than the prepending counterpart on a 7B LLM (Llama 2) in the runtime test.
[…]
https://github.com/BIT-NLP-GROUP/FlashBack
@@ 1 Introduction
Figure 1: A series of inference tests with simulated input and retrieved content on OPT and Llama 2 for comparing Appending Pattern with Prepending Pattern. (Maximum sequence length is denoted under the model name)
Figure 2: An illustration of the FlashBack pipelines. The left side of the diagram is the fine-tuning phase of our model with Marking Token and LoRA module. The inference procedure is on the right.
Large language models (LLMs) based on the Transformer architecture Vaswani et al. (2023) such as GPT, Llama, and OPT, etc. Brown et al. (2020); Touvron et al. (2023); Zhang et al. (2022) require enormous computational resources to keep their knowledge up-to-date Meng et al. (2023). Thus, Retrieval-Augmented Language Modeling (RALM) has emerged as a popular approach, enabling content generation that leverages external corpora to extend beyond the knowledge inherent in the model’s parameters, thereby reducing the computational cost of capturing up-to-date knowledge. The effectiveness of using RALM to improve the performance of LLM has been demonstrated by previous studies Guu et al. (2020); Lewis et al. (2021); Izacard et al. (2022); Wang et al. (2023b).
Recent works utilizing retrieved content simply by prepending retrieved contents to the input without updating language models exhibit the impressive potential of adapting off-the-shelf LLMs to the retrieved content Shi et al. (2023); Ram et al. (2023). Other studies build RALM through the pre-training of LLM Borgeaud et al. (2022), or engage in the continual pre-training of existing RALMs Izacard et al. (2022); Wang et al. (2023a).
However, these works are introduced with limitations. First, the off-the-shelf LLMs are not inherently trained to incorporate retrieved content, and extensive pre-training of LLMs for building RALM incurs high computational costs Lin et al. (2023). Second, although in-context methods have been effectively applied on off-the-shelf LLMs Ram et al. (2023), our work verified that prepending long input with In-Context RALM poses a high computational cost, which increases the inference time. To be specific, retrieved content changes with every retrieval stride tokens, and the computed key-value (KV) cache of the entire context is discarded, requiring re-computation for each change. The text generation process with a long input context can substantially increase the inference runtime, and the attention KV cache grows even larger on LLM with hundreds of billions of parameters Pope et al. (2022). Under these circumstances, recomputation of the KV cache becomes a computationally intensive task, entailing significant costs for RALM with continuous retrieval Ram et al. (2023).
@@
In this paper, we propose FlashBack to overcome these issues we mentioned above, a method that utilizes off-the-shelf LLMs efficiently without further pretraining of LLMs as shown in Figure 2, to address the inefficiency problem by breaking the dependence of the input to retrieved content in the context. To be specific, instead of prepending the retrieved content directly on the input, we utilize the Appending Context Pattern, in which the retrieved content is appended at the end of the input (see Figure 3). In our method, the KV cache of each input is static during inference, and the recomputation of the KV cache takes place only in the part of the retrieved content. Although there are advantages in applying this context pattern, it also presents an issue that degrades the performance of the RALM. Our appending pattern breaks the semantic coherence of the context, leading to a loss in perplexity when applying it to existing RALM (see Table 1). To solve this issue, inspired by Ren et al. (2023), FlashBack introduces a tunable Marking Token to adapt to our appending context pattern for restoring comparable performance.
@@
In particular, we use LoRA Hu et al. (2021) to fine-tune the Marking Token while keeping both the retriever and the LLM frozen. Therefore, FlashBack is an adaptable approach that can cooperate with other RALM that perform retrieval for every $n$ token, thereby aiming for efficient fine-tuning, inference and context-model alignment.
@@
  • Our work analyzed the prepending context in the FLOPs analysis and experiments. We find it is inefficient in RALM with long input due to redundant re-computation of the KV cache (see Figure 3 and section 3.3).
  • We find that the Appending Context Pattern can improve the inference efficiency of RALM. (Up to 4×4$\times$4 × faster in 7B Llama 2 as shown in Figure 1)
  • We exploit Marking Token that can be used for adapting the LLMs to new context patterns by efficient fine-tuning with LoRA instead of pre-training.
@@ 2 Background and Related Work
Retrieve-Read RALM
Previous works (Borgeaud et al., 2022; Lin et al., 2023; Ram et al., 2023; Shi et al., 2023) have created distinct modules for document selection and document reading, and this methodology can be defined as the Retrieve-Read RALM. In the recent RALM framework, particularly for those employing LLMs, the imperative is to align the retrieved documents with the specific requirements of the LLMs (Gao et al., 2024). AAR, REPLUG, and UPRISE are fine-tuning retrievers with frozen LLMs to achieve alignment (Cheng et al., 2023; Shi et al., 2023; Yu et al., 2023). RETRO uses frozen retrievers to build their own RALM (Borgeaud et al., 2022). In-Context RALM uses a frozen retriever for document selection and a frozen LLM for document reading without undergoing additional training for the LLM or the retriever (Ram et al., 2023). RAVEN (Huang et al., 2023) continually pre-train ATLAS (Izacard et al., 2022) for improving its in-context capabilities, which aligns the model to specific prompting strategy, and they also proposed Fusion-in-Context Learning for incorporating more in-context examples during its inference process.
Hence, our modeling is close to the RALM genre that uses decoder-only LLMs (Auto-regressive Models) and Retrieval-Read RALM that uses distinct modules for document selection and document reading (Borgeaud et al., 2022; Ram et al., 2023; Lin et al., 2023; Shi et al., 2023). Furthermore, our proposed method is classified as Input augmentation by recent study (Asai et al., 2024). After all, this type of RALM has recently been defined as Modular RAG (Retrieval-Augmented Generation) and is increasingly becoming the dominant norm in the RAG domain (Gao et al., 2024). Furthermore, it is worth mentioning that FinanceBench (Islam et al., 2023) also investigates the effectiveness of the context pattern. However, their efforts in modifying context positions are solely focused on inference accuracy rather than efficiency. GRIT-LM (Muennighoff et al., 2024) improves model inference efficiency by reusing embedding-based representations that only apply to the embedding model. Our work FlashBack, is a text-based method that is not limited to embedding models.
@@ 3.1 RALM with In-Context-Learning
For the trade-off between runtime cost and model performance, the retrieval stride $s$ and query length $\ell$ are proposed to perform the retriever call less frequently(at least, every s > 1 token). Therefore, the retrieval would be performed for every $s$ token, using the last $\ell$ tokens as designated retrieved content for input.
@@
$p(x_{1},...,x_{n})=\\ \prod_{j=0}^{n_{s}-1}\prod_{i=1}^{s}p_{\theta}(x_{s\cdot j+i}|[\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)});x_{<s\cdot j+i}]),$ (2)
@@
Our work followed this retrieval mechanism in addition to using different context patterns because this mechanism uses distinct retrieval modules and readers (LLMs) so that they can be optimized independently.
@@ 3.2 Context Pattern
In decoder-only transformer-based models, the computation of attention modules is related to the query of the current token and the key-value representations of preceding tokens (Vaswani et al., 2023). During inference, the key-value of the current token would be saved in the KV cache, eliminating the requirement in the subsequent step. We find that prepending retrieved content to the input has been prevalent in previous methods (Ram et al., 2023; Shi et al., 2023). However, the key-value computed for previous tokens becomes obsolete, and a re-computation of the new key-value is required for each retrieval step since the retrieved content varies with each retrieval.
To avoid the extensive cost of re-computation, we remove the autoregressive dependency between retrieved content and input in context by using an appending context pattern. Specifically, we append retrieved content to the input (at the end of the input) as shown in Figure 3.
@@
$p(x_{1},...,x_{n})=\\ \prod_{j=0}^{n_{s}-1}\prod_{i=1}^{s}p_{\theta}(x_{s\cdot j+i}|[x_{<s\cdot j+i};\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)})]).$ (3)
In this case, for each input, the change of retrieved content $\mathcal{R}_{\mathcal{C}}(q_{j}^{(s,\ell)})$ would not require a re-computation of new value, therefore preserving the computed value of previous tokens in KV cache $x_{<s\cdot j+i}$. (Algorithm pseudo code in Appendix A.)
@@ 3.4 Marking Token and Fine-tuning Choice
Since LLMs are not aligned explicitly with our appending pattern, we use the Marking Token and Low-Rank Adaption (LoRA) (Hu et al., 2021) to adapt them to the appending pattern while keeping the origin model weights frozen so that alignment is achieved without modifying the inherent ability of the LLMs. Marking Tokens are represented as two special prompt tokens denoting <MARK_L>,<MARK_R> which are added into the vocabulary of the model, and they are used for marking the boundary of the retrieved content in context (see Figure 2). Then, we fine-tune the model using LoRA with Marking Tokens applied for adapting the appending pattern. LoRA is a parameter-efficient fine-tuning (PEFT) for fine-tuning a subset of parameters in the pre-trained model while targeting for obtaining performance comparable to full model fine-tuning. Recent studies suggest that forgetting is inevitable in fine-tuning, but the PEFT method can facilitate less forgetting during the fine-tuning process, which does not greatly damage the performance of the pre-trained LLM (Kalajdzievski, 2024). In FlashBack, we opt for LoRA to demonstrate how fine-tuning a relatively minimal set of model weights enables our model to adapt to the appending context pattern while keeping pre-training model weights frozen. Our experiments show that FlashBack has significantly increased inference speed and can still maintain comparable perplexity (PPL) after fine-tuning. During fine-tuning, all parameters are frozen except for the embedding of <MARK_L>, <MARL_R>, and LoRA modules applied to attention layers.
@@ 3.5 FLOPs analysis of Marking Token and LoRA
If the rank of LoRA weight matrices is $r$, then the dimensions of the projection matrix are $[h,r]$. The average length of the retrieved content is $d$. We only equip LoRA modules, two projection matrices, for Key and Value matrices. Overall, the FLOPs of LoRA modules in appending context pattern can be formulated as: [^2] We only consider multiplication operations and omit some addition operations.
$C_{1}=2l(4bThr+bTh)\\ +2l\sum_{i=1}^{T/s}\left[4b(d+s)hr+b(d+s)h\right]\\ =\frac{2l(4r+1)bhT(d+2s)}{s}$ (5)
@@
$C_{\text{decreament}}=C_{0}-C_{1}\\ =\frac{2lTbh\left[(T+s)h-(4r+1)(d+2s)\right]}{s}$ (6)
@@ 3.6 Retriever
Our runtime test used a sparse model, BM25 (Robertson and Zaragoza, 2009), and the QA task is tested with the DPR retriever (Karpukhin et al., 2020) to demonstrate our idea. In practice, FlashBack intrinsically supports switching to other retrievers in a plug-and-play manner.
@@ 4.1 Run-time Improvement
We test OPT, GPT-2, and Llama2 in different model sizes and employ simulated retrieved content and input in inference speed tests. Our experiment included tests with varying input lengths. In addition, it involved examining the inference time under two distinct context patterns.